OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of sixty-six inches What is its diameter? Use twenty-two sevenths for pi

OpenStudy (ipwnbunnies):

\[Circumference = \pi*diameter\] Therefore: \[diameter = \frac{circumference}{\pi}\]

OpenStudy (anonymous):

66/3.14

OpenStudy (anonymous):

21

OpenStudy (anonymous):

thanks U guys rock